OpenStudy (anonymous):

Simplify 6(A + 2) + 5A. 23A 12A + 11 11A + 12 Simplify 7(x + 2) + 12. 7x + 14 7x + 21 7x + 26 Simplify 18 - 2[x + (x - 5)]. 13 - x 13 - 4x 8 - 4x 28 - 4x Simplify 8(A + 2B) + 6(2A + B). 20A + 22B 20A + 3B 20A + 2B i already did it i just want to make sure they are right, better safe than sorry. thank you :)

OpenStudy (sleepyjess):

What did you choose for each one? :)

OpenStudy (anonymous):

1) A 2) B 3) D 4)B

OpenStudy (sleepyjess):

Only 3 is correct, can you show me your work for the others?

OpenStudy (anonymous):

all i did was distirbute and multiply, i probably multiplied wrong

OpenStudy (sleepyjess):

Ok, for the first one, we have 6(A + 2) + 5A, which is 6A + 6(2) + 5A

OpenStudy (anonymous):

1)a. 6(a+2)+5a =6a+12+5a =11a+12 the answer is (11a +12)

OpenStudy (anonymous):

okay thank you both, how do i solve the other ones?

OpenStudy (sleepyjess):

For the second one, 7x + 7(2) + 12. What is 7*2?

OpenStudy (anonymous):

14

OpenStudy (sleepyjess):

Yep, so 7x + 14 + 12. 14 + 12?

OpenStudy (anonymous):

26

OpenStudy (sleepyjess):

yep, so the final answer is 7x + 26

OpenStudy (anonymous):

b. the answer is 7x+26

OpenStudy (anonymous):

thank you, can you help me with the last one also?

OpenStudy (sleepyjess):

Yep, 8(A + 2B) + 6(2A + B) is the same as 8A + 8(2B) + 6A + 6B

OpenStudy (sleepyjess):

What is 8*2B?

OpenStudy (anonymous):

16B

OpenStudy (sleepyjess):

Yep, now rearrange it so like terms are together, 8A + 6(2A) + 16B + 6B. 6*2A = ?

OpenStudy (anonymous):

c.the answer is -4x2+18+20x

OpenStudy (anonymous):

12A

OpenStudy (sleepyjess):

@mizanin , Please do not give out direct answers. Welcome to OpenStudy!

OpenStudy (sleepyjess):

Yep, so 8A + 12A + 16B + 6B 8A + 12A

OpenStudy (anonymous):

d.10a+22b

OpenStudy (anonymous):

please stop, thats not helping me learn @mizanin

OpenStudy (sleepyjess):

Now we have 8A + 12A + 16B + 6B. What is 8A + 12A?

OpenStudy (anonymous):

20A

OpenStudy (sleepyjess):

yes, now 20A + 16B + 6B. What is 16B + 6b?

OpenStudy (anonymous):

22B

OpenStudy (sleepyjess):

Yep, so the final answer is 20A + 22B :)

OpenStudy (anonymous):

thank you so much for helping me! :)

OpenStudy (sleepyjess):

Thank you for participating and trying to understand the material instead of taking the (incorrect) answers that @mizanin was giving :)

OpenStudy (anonymous):

ohh sorry its 8(a+2b)+6(2a+b) 8a+16b+12a+6b =22a+22b so the answer is 22a+22b

OpenStudy (anonymous):

and am really sorry its just that i was being disturbed by a child guess u forgive